Operations Report – February 14th

Crew 221 Operations Report 14Feb2020

SOL: 12

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: on (at 7:16 pm) 89%

Hours run: 10:40

From what time last night: 8:20pm

To what time this morning: 7:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 73%

Diesel Reading – 73%

Station Propane Reading – 47%

Ethanol Free Gasoline – N/A

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 145 gals

Static to Loft Pump used – yes

Water in GreenHab – 157.5 gals used and no water left in the greenhab at this moment, please refer to greenhab report.

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes

Hours: 142.8

Beginning charge: 100%

Ending charge: 47%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 131.9

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 53%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: no

Hours: 137.2

Beginning charge: 100%

Ending charge:

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? not used

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment finished, watering of the plants carried out as per procedure.

Summary of ScienceDome operations: one oven is operating in the lab, the incubator is non-functional, microbiology experiments going on and cultures are being grown in agar plates, all activities set to end by tomorrow.

Summary of RAM operations: RAM cleaned

Summary of any observatory issues: nominal

Summary of health and safety issues: nothing to report.

Questions, concerns and requests to Mission Support: Cleaning of habitats, food inventory carried out.

Copyright © The Mars Society. All rights reserved. | Main Site